OpenStudy (anonymous):

find the derivative of the trigonometric function… h(theta)=5 theta sec theta+theta tan theta CALCULUS HELP PLEASE!

OpenStudy (anonymous):

please help me!!!!!!!!!!!

OpenStudy (anonymous):

Alright, so you have: \[h(\theta)=5\theta\frac{1}{cos(\theta)}+\theta tan(\theta)\] And you want to find \(h'(\theta)\) RIght?

OpenStudy (candy13106):

keith plz help man on my question a fewer more and than i got to go tot bed

OpenStudy (anonymous):

@KeithAfasCalcLover yes!!!

OpenStudy (candy13106):

plz some one help i need help plz

OpenStudy (anonymous):

group theta like this:\[h(\theta) = \theta*(5*\sec(\theta)+\tan(\theta)\]

OpenStudy (anonymous):

\[h(\theta)=\theta\left(\frac{5}{cos(\theta)}+tan(\theta)\right)\] And use the product rule: \[h'(\theta)=\theta\left(\frac{5}{cos(\theta)}+tan(\theta)\right)'+\left(\frac{5}{cos(\theta)}+tan(\theta)\right)\] \[h'(\theta)=\theta\left(5\left(\frac{1}{cos(\theta)}\right)'+\frac{1}{cos^2(\theta)}\right)+\left(\frac{5}{cos(\theta)}+tan(\theta)\right)\]

OpenStudy (anonymous):

\[h'(\theta)=\theta\left(5\left(\frac{-1}{cos^2(\theta)}\times-sin(\theta)\right)+\frac{1}{cos^2(\theta)}\right)+\left(\frac{5}{cos(\theta)}+tan(\theta)\right)\] \[h'(\theta)=\theta\left(\frac{5sin(\theta)+1}{cos^2(\theta)}\right)+\left(\frac{5}{cos(\theta)}+tan(\theta)\right)\]

OpenStudy (anonymous):

\[h'(\theta)=\frac{\theta[5sin(\theta)+1]}{cos^2(\theta)}+\frac{5+sin(\theta)}{cos(\theta)}\]

OpenStudy (anonymous):

I probably could go a lil bit further but ill stop there. Cool Hannah?

OpenStudy (anonymous):

@KeithAfasCalcLover omggg thank you so muchh!!!!!! you are the best

OpenStudy (anonymous):

Haha I try ;-) Im glad you liked it. Anytime.

OpenStudy (anonymous):

@KeithAfasCalcLover couldi ask you some more questions later on? ive been having so much trouble doing this stupid hw:'(

OpenStudy (anonymous):

Yeah sure! Whenever i'm on i'd love to answer some Calc questions! HAVE YOU SEEN MY NAME? Haha

OpenStudy (anonymous):

@KeithAfasCalcLover hahaha omg totally didnt notice that until now;) i think i really like that name <333

OpenStudy (anonymous):

LOL Well i-i-im flattered hahaha

OpenStudy (anonymous):

What Calculus do you have? High School? University?

OpenStudy (anonymous):

@KeithAfasCalcLover haha;) im taking calculus ab online so i guess it's highschool calc:)

OpenStudy (anonymous):

Ahh I see. Just out of curiosity, why online? If you don't mind me asking :p

OpenStudy (anonymous):

@KeithAfasCalcLover totally dont mind! im taking bc at school and i thought taking it ab online will help me to review stuff twice and boost up my gpa...but now i kinda regret taking it with all these overwhelming problems:O

OpenStudy (anonymous):

Haha naww don't worry. Ehh whats bc calculus and ab calculus? Im not familiar with the terminology haha

OpenStudy (anonymous):

@KeithAfasCalcLover they are just highschool calc and bc is a little harder than ab and covers more topics but they are basically the same thing:p

OpenStudy (anonymous):

Lol ahh I see. So how far did you guys get so far?

OpenStudy (anonymous):

Probably by now, covered limits, and derivatives right?

OpenStudy (anonymous):

@KeithAfasCalcLover yea!! i think we are on derivative

OpenStudy (anonymous):

@KeithAfasCalcLover determine the point(s) at which the graph of the function has a horizontal tangent line. f(x)=x^2/(x^2+1) how do you do this one?:'(

OpenStudy (anonymous):

haha its not that bad. The "horizontal line" means that the line has a slope of zero right?

OpenStudy (anonymous):

@KeithAfasCalcLover yes!!

OpenStudy (anonymous):

So then the "horizontal tangent line" can be found at the x-values where the derivative is zero since the derivative describes the instantaneous slopes of the function at any point! So you have : \[f(x)=\frac{x^2}{x^2+1}\] You can find the derivative by applying the quotient rule which says: \[f'(x)=\left(\frac{g(x)}{h(x)}\right)'=\frac{g'(x)h(x)-h'(x)g(x)}{[h(x)]^2}\] So lets break it down: What is our g(x) and h(x)?

OpenStudy (anonymous):

@KeithAfasCalcLover g(x) is x?? im not suree

OpenStudy (anonymous):

;(

OpenStudy (anonymous):

Lol aww don't cry... \[g(x)=numerator\] \[h(x)=denominator\]

OpenStudy (anonymous):

Little bit better?

OpenStudy (anonymous):

So therefore \(g(x)=x^2\) and \(h(x)\) would be?

OpenStudy (anonymous):

@KeithAfasCalcLover will it be x^2+1??

OpenStudy (anonymous):

Yeah! So now lets take the derivative of \(g(x)\) and \(h(x)\). Ill do the first and you do the second deal? :-)

OpenStudy (anonymous):

\[g'(x)=(x^2)'=2x\]

OpenStudy (anonymous):

@KeithAfasCalcLover yay!!! ok deal!;)

OpenStudy (anonymous):

Haha done! Your turn

OpenStudy (anonymous):

h'(x)=2x..........?

OpenStudy (anonymous):

@KeithAfasCalcLover did i get it???

OpenStudy (anonymous):

NICE haha

OpenStudy (anonymous):

@KeithAfasCalcLover wootwooot:D

OpenStudy (anonymous):

So we know: \(g(x)=x^2\) \(g'(x)=2x\) \(h(x)=x^2+1\) \(h'(x)=2x\) And we know that: \[f'(x)=\left(\frac{g(x)}{h(x)}\right)'=\frac{g'(x)h(x)-h'(x)g(x)}{[h(x)]^2}\] So lets plug everything in!: \[f'(x)=\left(\frac{x^2}{x^2+1}\right)'=\frac{2x(x^2+1)-2x(x^2)}{(x^2+1)^2}\]

OpenStudy (anonymous):

AAAAAND We can simplify a little: \[f'(x)=\frac{2x^3+2x-2x^3}{(x^2+1)^2}\] \[f'(x)=\frac{2x}{(x^2+1)^2}\]

OpenStudy (anonymous):

But we're still not done. We need to find out when \(f'(x)=0\)

OpenStudy (anonymous):

So lets set it to zero: \[0=\frac{2x}{(x^2+1)^2}\] \[0(x^2+1)^2=2x\] \[0=2x\] \[0/2=x\] \[x=0\] So when will there be a horizontal tangent line? :-)

OpenStudy (anonymous):

when x=0????

OpenStudy (anonymous):

YEAH. Which makes sense! Because:

OpenStudy (anonymous):

The tangent line at \(x=0\) would be horizontal, you can see!

OpenStudy (anonymous):

@KeithAfasCalcLover omggg!! yes i get it!! THANKS SO MUCHH

OpenStudy (anonymous):

Haha great! Anytime!

OpenStudy (anonymous):

Yeahhh. Im in grade 12 now! Ill be getting the first introduction of calculus in like March haha

OpenStudy (anonymous):

@KeithAfasCalcLover omg you go to highschool? how are you soo good at math?!?

OpenStudy (anonymous):

Lol haha Well I have a problem. If I have a question about something or something I need to know, I find it out haha. So I heard of calculus in grade 11 and heard it's applications. So I got curious and looked at derivatives and limits. Then I heard of the idea of area under a curve and found that it was related to the integral so I learned that. Then I learned ways to apply it to multivariable approaches so like multivariable calculus. And that's where im at right now. I always had a thing for math. It comes to me like English comes to people who speak. Its really like a language haha. yeahhhh...